# Hard constraint learning approaches with trainable influence functions for evolutionary equations

### v2

@@ Abstract
Deep learning , Physics-informed neural networks , Hard constraint , Evolutionary equation
@@ 4.2 Allen-Cahn equation
0.5(Training) ${\bf 1.0372\times 10^{-3}}$ ${\bf 3.5572\times 10^{-4}}$ ${\bf 5.2209\times 10^{-3}}$
@@ 4.3 Korteweg-de Vries equation
4.4 A 3D heat conduction problem with large gradients
In practical applications such as materials science [25], it is often necessary to investigate the thermal stability and durability of materials when the temperature changes drastically in a short time. In this section, we design a 3D space model to test the effectiveness of our method for this type of problem.
Consider the following three-dimensional heat equation,
$\begin{cases}u_{t}=u_{xx}+u_{yy}+u_{zz}+f(x,y,z,t),&(x,y,z)\in(-1,1)^{3},t\in(0,1],\\ u(x,y,z,0)=(1.0+0.1\cdot x\cdot y\cdot z)\cdot e^{\frac{1}{1+\varepsilon}},&(x,y,z)\in(-1,1)^{3},\\ u(\pm 1,y,z,t)=(1.0\pm 0.1\cdot y\cdot z)\cdot e^{\frac{1}{(t-1)^{2}+\varepsilon}},&(y,z)\in(-1,1)^{2},t\in(0,1],\\ u(x,\pm 1,z,t)=(1.0\pm 0.1\cdot x\cdot z)\cdot e^{\frac{1}{(t-1)^{2}+\varepsilon}},&(x,z)\in(-1,1)^{2},t\in(0,1],\\ u(x,y,\pm 1,t)=(1.0\pm 0.1\cdot x\cdot y)\cdot e^{\frac{1}{(t-1)^{2}+\varepsilon}},&(x,y)\in(-1,1)^{2},t\in(0,1].\\ \end{cases}$ (4.7)
Its exact solution is given by
$u(x,y,z,t)=(1.0+0.1\cdot x\cdot y\cdot z)\cdot e^{\frac{1}{(t-1)^{2}+\varepsilon}}.$ (4.8)
where $\varepsilon$ is a positive constant, and here we take $\varepsilon=0.15$. $f(x,y,z,t)$ is derived from the exact solution.
In this model, the equation varies within $[0.9,1.1]$ in the spatial domain. Consequently, its range is primarily determined by the parameter $\varepsilon$, which displays well the performance of THC-PINNs respond to challenge of abrupt temporal variations. When $\varepsilon$ is large, $u$ varies smoothly across the domain; when $\varepsilon$ is small, $u$ exhibits sharp changes near $t=1$, highlighting the multi-scale nature of the problem. Thus, we employ the THC-PINN method to solve the model with $\varepsilon=0.15$ and compare its result with the conventional PINN method. Regarding temporal domain partitioning, we adpot adaptive temporal domain partitioning algorithm in Sec.3. Table 4.11 lists the approximate $L_{2}$ relative error $\mathcal{D}$ used for adaptive partitioning. Based on this data, we solve the heat equation using $4$ intervals.
@@ 4.4 A 3D heat conduction problem with large gradients
Table 4.11: Adaptive time interval partitioning for the heat equation.
@@
Table 4.12: Solving the heat equation with different PINN method.
Table 4.12 lists the training results obtained using the conventional PINN method and the THC-PINN method with four intervals. Figure 4.12 shows the comparison between the THC-PINN solution and the exact solution at $t=1$. The predictive solution obtained by our method closely approximates the reference solution. Figure 4.13 illustrates the THC-PINN solution at $z=1$ for $t=0.5,0.6,\cdots,1.0$ with variation being $[10.9904,866.5444]$. Our method copes well with multi-scale characteristics in time. Notably, the same as the AC equation, the heat equation is also a parabolic equation, so the PINN solution to Eq. (4.7) exhibits excellent extrapolation capabilities, with $p_{i}$ trained up to the right endpoint of the intervals.
@@
Figure 4.13: THC-PINN solution at $z=1,t=0.5,0.6,\cdots,1.0$.

### v3

@@ Abstract
Deep learning , Physics-informed neural networks , Hard constraint , Evolutionary equations
@@ 4.2 Allen-Cahn equation
0.50(Training) ${\bf 1.0372\times 10^{-3}}$ ${\bf 3.5572\times 10^{-4}}$ ${\bf 5.2209\times 10^{-3}}$
@@ 4.3 Korteweg-de Vries equation
4.4 A three-dimensional heat conduction problem with large temporal gradients
In materials science [25], it is often necessary to study the thermal stability and durability of materials when the temperature changes dramatically in a short period of time. Here we test the effectiveness of our method for the problem with large temporal gradients by
$\begin{cases}u_{t}=\Delta u+f(\bm{x},t),&\bm{x}\in\Omega,t\in(0,1],\\ u(\bm{x},t)=g_{0}(\bm{x}),&t=0,\\ u(\bm{x},t)=g_{D}(\bm{x},t),&\bm{x}\in\partial\Omega,\end{cases}$
here $\Omega=(-1,1)^{3}$, $g_{D}$, $g_{0}$ and $f$ are derived by the exact solution
$u(x,y,z,t)=(1.0+0.1xyz)e^{\frac{1}{(t-1)^{2}+\varepsilon}},$
where $\varepsilon$ is a positive constant. It can be seen that the variation of solutions in the spatiotemporal domain mainly depends on the size of $\varepsilon$ in the time term. If $\varepsilon$ is large, the variation of $u$ in the entire computational domain will be smooth; If $\varepsilon$ is very small, the value of $u$ will sharply change around $t=1$, reflecting the multi-scale characteristics in the time direction.
Here we employ the proposed method THC-PINN to solve this problem with $\varepsilon=0.15$. We adopt the adaptive temporal domain partitioning algorithm in Section 3 to divide temporal domain. Table 4.11 lists the approximate $L_{2}$ relative error $\mathcal{D}$ used for adaptive partitioning. Based on this data, we solve the three-dimensional heat conduction equation using $4$ intervals. Notably, the same as the AC equation, the heat conduction equation also belongs to the parabolic type, and the solution has strong extrapolation ability. $p_{i}$ is trained to the right endpoint of the intervals.
@@ 4.4 A three-dimensional heat conduction problem with large temporal gradients
Table 4.11: Adaptive time interval partitioning for the 3D problem.
@@
Table 4.12: The computational errors with different PINN methods.
Table 4.12 presents the computational errors of standard PINNs and THC-PINNs with four intervals. Figure 4.12 depicts the contours of exact solution, standard PINN solution, THC-PINN solution at $t=1$. Figure 4.13 shows the slice of the THC-PINN solution on the face $x=y$ at time $t=0.90,0.95,1.00$. These results show the superiority of our method for the time multi-scale problem.
@@
Figure 4.13: THC-PINN solution at $x=y$ with different time.